OpenStudy (anonymous):

gustavo weights 202 pounds he wants to weigh 188 pounds in 4 months what is the average number of pounds he should plan to lose each month to reach his goal

OpenStudy (binary3i):

he is 202 pounds and he want to become of 188 pounds. he wants to loos 202-188= 14 pounds in 4 months so each month lose is 14/4 =7/2 pounds

OpenStudy (anonymous):

thanks

OpenStudy (binary3i):

welcome

OpenStudy (anonymous):

could you help with 1 more? emily is constructing a triangle one side is 3 inches long and another is 5 inches long which of the following is the smallest measure emily could use for the third side of the triangle

OpenStudy (binary3i):

i dont know actually.